OpenStudy (anonymous):

Confusion! Let f(x)=x^2+5x, g(x)=6-2x, h(x)=2/x-3; find the following value. f(6)-4=

OpenStudy (anonymous):

f(6) -4=((6)^2+5(6))-4= 36+30-4= 62

OpenStudy (anonymous):

Whoa thanks! Can you please tell me how you got that?

OpenStudy (anonymous):

Oh wait, just plug in?

OpenStudy (anonymous):

Yeah f(x) is x^2+5x. So f(6) is plugged into the x giving (6)^2+5(6)= 36+30=66. The problem asks for f(6)-4 so that is 66-4=62

OpenStudy (anonymous):

Yeah

OpenStudy (phi):

people don't like to type so they abbreviate with things like f(x) it means go look for the definition, and replace x with whatever number I put in example f(2) means plug in 2 for x. saves typing.

OpenStudy (anonymous):

Oh I see. But I tried this one out: Let f(x)=x^2+5x, g(x)=6-2x, h(x)=2/x-3; find the following value. f(-3). I got -24, is this correct?

OpenStudy (phi):

9-15= -6

OpenStudy (phi):

-3*-3 is +9

OpenStudy (anonymous):

(-3)^2+5(-3)= 9-15=-6

OpenStudy (anonymous):

OH o.k. Now I understand. Thanks guys! :)